# What’s Mine becomes Yours: Defining, Annotating and Detecting Context-Dependent Paraphrases in News Interview Dialogs

### v1

@@ 5.1 Annotation Task
5.2 Plausible Label Variation
The task of annotating context-independent paraphrases is already difficult. Disagreements between human annotators are common (Dolan and Brockett, 2005; Krishna et al., 2020; Kanerva et al., 2023) — even with extensive manuals for annotators (Kanerva et al., 2023). In related semantic tasks like textual entailment,888 Paraphrase classification has been repeatedly equated to (bi-)directional entailment classification (Dolan and Brockett, 2005; Androutsopoulos and Malakasiotis, 2010) disagreements have been linked to plausible label variations inherent to the task (Pavlick and Kwiatkowski, 2019; Nie et al., 2020; Jiang and de Marneffe, 2022).
5.3 Annotator Training
When annotating paraphrases, the instructions for annotators are often short, do not explain challenges and rely on annotator intuitions (Dolan and Brockett, 2005; Lan et al., 2017).999 For example, instructions are to rate if two sentences “mean the same thing” (Dolan and Brockett, 2005) or are “semantically equivalent” (Lan et al., 2017). In contrast, Kanerva et al. (2023) recently used an elaborate 17-page manual. However, they relied on only 6 expert annotators that might not be able to represent the full complexity of the task (§5.2). We aim for a trade-off between short intuition-based and long complex instructions that facilitates recruitment of a larger number of annotators: an accessible example-centric, hands-on 15-minute training of annotators that teaches our operationalization of context-dependent paraphrases (§3). We provide (1) a short paraphrase definition, (2) examples of context-dependent paraphrases showing clear and approximate equivalence (c.f. Table 2), (3) examples of common difficulties with paraphrase classification in dialog (c.f. Table 3 and §3), and use (4) a hands-on approach where annotators have to already classify and highlight paraphrases after receiving instructions. Only once they make the right choice on what is (Table 2) and is not a paraphrase (Table 3) and highlight the correct spans they are shown the next set of instructions. Only annotators that undergo the full training and pass two comprehension and two attention checks are part of our released dataset. Overall, 49% of the annotators who finished the training passed it. See App. C for the instructions and further details.
@@ 5.3 Annotator Training
llama 2 7B 1% 0.66 0.49 0.98 52% 0.28 0.44
vicuna 7B 0% 0.29 0.67 0.19 26% 0.31 0.46
Mistral 7B Instruct v0.2 3% 0.62 0.66 0.58 54% 0.41 0.49
openchat 3.5 0% 0.66 0.76 0.58 55% 0.40 0.55
gemma 7B 1% 0.64 0.66 0.63 41% 0.24 0.49
Mixtral 8x7B Instruct v0.1 0% 0.74 0.73 0.74 57% 0.40 0.54
Llama 2 70B 0% 0.66 0.72 0.61 69% 0.27 0.56
GPT-4 0% 0.81 0.78 0.84 9% 0.63 0.69
DeBERTa v3 large - 0.66 0.82 0.56 - 0.45 0.64
Table 7: Modeling Results. We boldface the best and underline the second best performance. We display the extraction error of predictions from generative models and, for classification, the F1, precision and recall score as well as, for highlights, the Jaccard Index for the guest and host utterances. Higher values are better ($\uparrow$) except for extraction errors ($\downarrow$). GPT-4 is the best classification model, while, overall, DeBERTa is the best highlight model.
@@ 5.5 Results
We discuss annotations results (tables 1, 4, 5) on our datasets BALANCED, RANDOM and PARA.
Classification agreement as an indicator of variation. Agreement for classification is relatively low (Table 1). We inspect a sample of 100 annotations on the RANDOM set and manually assess annotation quality. 90% of the annotations can be said to be at least plausible (see Table 6 for low quality and Table 1 for plausible variation examples), which is in line with the fact that we only use high quality annotators (§5.3). As a result, we assume that most disagreements are due to the inherent plausible label variation of the task (§5.2).
Higher agreement on paraphrase position. Krippendorff’s $\alpha$ on the highlights is higher than in other areas101010 E.g., 0.41 for hate speech (Carton et al., 2018) or 0.35 for sentiment analysis (Sullivan Jr. et al., 2022). Because of the different tasks these values are not exactly comparable. (see Table 5). We also calculate the “Intersection-over-union” between the highlighted words (i.e., Jaccard Index), a common and interpretable evaluation measure for annotator highlights (Mendez Guzman et al., 2022; Mathew et al., 2021; Malik et al., 2021). It seems that while annotations vary on whether there is a paraphrase or not, they agree frequently on the position of the possible paraphrase. On average, at least 50% of the highlighted words are the same between annotations.111111 100% overlap in highlighting is uncommon. DeYoung et al. (2020) consider two highlights a match if Jaccard is greater than 50%. Agreement is higher on the host utterance, because on average the host utterance is shorter than the guest utterance (33 $<$ 85 words).
@@ 6 Modeling
Token Classifier. Similar to Wang et al. (2022a), we fine-tune a large DeBERTa model121212microsoft/deberta-v3-large (He et al., 2020) on token classification to highlight the paraphrase positions using the 3,896 training annotations (for hyperparameters, see App. D.2). We consider a model to have predicted a paraphrase for a pair if at least one token is highlighted in both texts. We average performances over three seeds.
In-Context Learning. We further prompt the following generative models (see URLs in App. D.1) to both classify and highlight the position of paraphrases: Llama 2 7B and 70B (Touvron et al., 2023), Vicuna 7B (Zheng et al., 2023), Mistral 7B Instruct v0.2 (Jiang et al., 2023), Openchat 3.5 (Wang et al., 2023), Gemma 7B (Team et al., 2024), Mixtral 8x7B Instruct v0.1 (Jiang et al., 2024) and GPT-4 (Achiam et al., 2023). We design the prompt to be as close as possible to the annotator training using a few-shot setup (Brown et al., 2020; Zhao et al., 2021) with all 8 examples shown during annotator training. We also provide explanations in the prompt (Wei et al., 2022b; Ye and Durrett, 2022) and use self-consistency by prompting the models 10 (GPT-4 and Llama 70B: 3) times (Wang et al., 2022b). For the prompt and further hyperparameter settings see App. D.1.
Split # (guest, host)-pairs # annotations
Train 420 3896
Dev 88 842
Test 92 843
@@
Overall, GPT-4 and Mixtral-8x7B achieve the best results in paraphrase classification. In highlighting, our DeBERTa token classifier and GPT-4 achieve the best overlap with human annotations. However, due to problems with extracting highlights from model responses (e.g., hallucinations, see App. D.3), our fine-tuned DeBERTa token classifier is probably the best choice to extract the position of paraphrases. We provide our best-performing DeBERTa model (model with seed 201 and F1 score of 0.72) on the Hugging Face Hub. [^13] LINK WILL BE ADDED AFTER BLIND REVIEW
Preds Shortened Examples
T
G
D
G: [Trump] is appointing very young judges. H: if you’re 50-plus, you’re probably too old for the Trump Administration to be seriously considered for judgeship.
Error Analysis. We consider the best-performing classification and highlighting models for error analysis, i.e., GPT-4 and DeBERTa. We manually analyze a sample of misclassifications, for examples see Table 9. Overall, the classification quality is better for GPT-4. The DeBERTa token classifier misses some more obvious paraphrases, while GPT-4 classifies some non-paraphrases. For both models, the items with incorrect predictions also show higher human disagreement. The average entropy for human classifications is lower for the correct (0.42 for DeBERTa, 0.45 for GPT-4) than for the incorrect model predictions (0.65 for DeBERTa, 0.67 for GPT-4). Highlights are of lower quality for GPT-4 (see App. D.3). DeBERTa generally highlights shorter spans of text, while GPT-4 usually highlights complete (sub-)sentences (e.g., “Yes, you’ve been doing this for a while now.” highlighted by GPT-4 vs. “you’ve … for a while” by DeBERTa in Figure 1).
@@ 7 Conclusion
A majority of work on paraphrases in NLP has looked at the semantic equivalence of sentence pairs in context-independent settings. However, the human dialog setting is highly contextual and typical methods fall short. We provide an operationalization of context-dependent paraphrases and an up-scalable hands-on training for annotators. We demonstrate the annotation approach by providing 5,581 annotations on a set of 600 turn pairs from news interviews. Next to paraphrase classifications, we also provide annotations for paraphrase positions in utterances. In-context learning and token classification both show promising results on our dataset. With this work, we hope to contribute to the automatic detection of paraphrases in dialog.
@@
Even though the number of our unique text pairs is relatively small, we release a high number of high quality annotations per text pair (5,581 annotations on 600 text pairs). Releasing more annotations on fewer “items” (here: text pairs), has increasingly been more common in NLP (Nie et al., 2020; Sap et al., 2022). Further, big datasets become less necessary with better generative models: Using only eight paraphrases pairs in our prompt already led to promising results. We further use the full 3,896 annotations from the training set to train a token classifier showing competitive results with the open generative models. However, the token classifier would probably profit from a bigger dataset.
@@
In this work we collect a high number of human annotations per item and highlight the plausible label variation in our dataset. However, we use hard instead of soft-evaluation approaches (Uma et al., 2021) for the computational models. We do this because, among others, extracting label distributions for in-context learning is challenging (Hu and Levy, 2023; Lee et al., 2023). We leave the development of a soft evaluation approach to future work.
@@
We confirm to have read and that we abide by the ACL Code of Ethics. Beside the mentioned ethical considerations, we do not forsee immediate risks of our work.
@@ B.1 Preprocessing
1. Filtering for 2-person interviews. We filter 49,420 NPR and 414,176 CNN interviews from Zhu et al. (2021) for 2-person interviews only. This can be challenging: In the speaker list, authors sometimes have non-unique identifiers (e.g., ‘STEVE PROFFITT’, ‘PROFFITT’ or ‘S. PROFFITT’ refer to the same speaker). If one author identifier string is contained in the other we assume them to be the same speaker.151515 There might be other cases where different string identifiers in the dataset refer to the same speaker although they are not substrings of the other (e.g., ‘S. PROFFITT’ and ‘STEVE PROFFITT’). For a randomly sampled selection of 44 interviews that were identified as more than 2 person interviews, 12 contained errors in the matching. 2/12 were the result of typos and 10/12 were the result of additions to the name like “(voice-over)” or “(on camera)”. We generally assume the first speaker to be the host. We remove 538 NPR and 1,917 CNN interviews because the identifier of the second speaker includes the keywords “host” or “anchor” — thus contradicting our assumption. This leaves 14,000 NPR and 50,301 CNN 2-person interviews.
@@
2. Removing first and last turns of an interview. The first turns in our 2-person interviews are usually (reactions to) welcoming addresses and acknowledgments by host and guest [^16] For example, “I’m Farai Chideya.” “Welcome.” “Thank you.” , while the last often contain goodbyes or acknowledgments171717 For example the last 3 turns in the considered NPR-4 interview: “Well, Dr. Hader. Thanks for the information.”, “Well, thank you for helping share that information […]”, “Well, thanks again. Dr. Shannon Hader […]”. We remove the first two and the last two (guest, host)-pairs. This step removes 2,409 NPR and 26,419 CNN interviews because they are fewer than 5-turns long. For the remaining interviews, this removes 34,773 NPR and 71,646 CNN (guest, host)-pairs.
@@
3. Removing short and long utterances. We further remove short guest utterances of 1–2 words as they leave not much to paraphrase.181818 We manually looked at a random sample of $0.3\%\approx 48$ such pairs. The 1-2 token guest utterances are mostly (40/48) assertions of reception by the guest (e.g., “Yes.”, “Exactly. Exactly.”, “That’s right”). Some are signals of protest (4/48) (e.g., “Hey, man.”, “Yes, but…”, “Hold on.”). None of them were reproduced by the host in the next turn. 3,540 NPR and 12,675 CNN pairs are removed like this. We also remove pairs where the host utterance consists of only 1–2 words.191919 We manually looked at a random sample of $0.3\%\approx 37$ such pairs. The 1–2 tokens host utterances are mostly (28/37) assertions of reception by the host (e.g., “Yeah.”, “Yes.”, “Sure.”, “Right.”, “Right. Right.”, “Ah, okay.”). Some are requests for elaboration (5/37) (e.g., “How so?”, “Like?”, “Four?”) or reactions (3/37) (e.g., “Wow!”, “Oh, interesting.”). Only one example “Four?” was reproducing content in the form of a repetition. . 2,940 NPR and 11,389 CNN pairs are removed like this. We also remove pairs where guest or host utterance consist of more than 200 words.202020 200 is the practical limit for the number of words for the chosen type of question (i.e., ‘Highlight” Question) in the used survey hosting platform (i.e., Qualtrics). It also limits annotation time per question. Overall, this leaves 148,522 (guest, host)-pairs in 34,419 interviews for potential annotation, see Table 10.

### v2

@@ 5.2 Plausible Label Variation
llama 2 7B 1% 0.66 0.49 0.98 59% 0.34 0.44
vicuna 7B 1% 0.29 0.67 0.19 32% 0.30 0.46
Mistral 7B Instruct v0.2 3% 0.62 0.66 0.58 66% 0.40 0.51
openchat 3.5 0% 0.66 0.76 0.58 64% 0.46 0.50
gemma 7B 1% 0.64 0.66 0.63 48% 0.24 0.51
Mixtral 8x7B Instruct v0.1 0% 0.74 0.73 0.74 65% 0.35 0.52
Llama 2 70B 0% 0.66 0.72 0.61 71% 0.29 0.56
GPT-4 0% 0.81 0.78 0.84 17% 0.67 0.71
DeBERTa v3 large AGGREGATED - 0.73 0.67 0.81 - 0.52 0.66
DeBERTa v3 large ALL - 0.66 0.82 0.56 - 0.45 0.64
Table 8: Modeling Results. We boldface the best and underline the second best performance. We display the extraction error of predictions from generative models and, for classification, the F1, precision and recall score as well as, for highlights, the Jaccard Index for the guest and host utterances. Higher values are better ($\uparrow$) except for extraction errors ($\downarrow$). GPT-4 is the best classification model, while, overall, DeBERTa is the best highlight model as it does not lead to any extraction errors.
5.3 Annotator Training
When annotating paraphrases, the instructions for annotators are often short, do not explain challenges and rely on annotator intuitions (Dolan and Brockett, 2005; Lan et al., 2017).121212 For example, instructions are to rate if two sentences “mean the same thing” (Dolan and Brockett, 2005) or are “semantically equivalent” (Lan et al., 2017). In contrast, Kanerva et al. (2023) recently used an elaborate 17-page manual. However, they relied on only 6 expert annotators that might not be able to represent the full complexity of the task (§5.2). We aim for a trade-off between short intuition-based and long complex instructions that facilitates recruitment of a larger number of annotators: an accessible example-centric, hands-on 15-minute training of annotators that teaches our operationalization of context-dependent paraphrases (§3). We provide (1) a short paraphrase definition, (2) examples of context-dependent paraphrases showing clear and approximate equivalence (c.f. Table 2), (3) examples of common difficulties with paraphrase classification in dialog (c.f. Table 3 and §3), and use (4) a hands-on approach where annotators have to already classify and highlight paraphrases after receiving instructions. Only once they make the right choice on what is (Table 2) and is not a paraphrase (Table 3) and highlight the correct spans they are shown the next set of instructions. Only annotators that undergo the full training and pass two comprehension and two attention checks are part of our released dataset. Overall, 49% of the annotators who finished the training passed it. See App. C for the instructions and further details.
@@ 5.5 Results
We discuss annotations results (tables 1, 4, 6) on our datasets BALANCED, RANDOM and PARA.
Classification agreement as an indicator of variation. Agreement for classification is relatively low (Table 1). We inspect a sample of 100 annotations on the RANDOM set and manually assess annotation quality. 90% of the annotations can be said to be at least plausible (see Table 7 for low quality and Table 1 for plausible variation examples), which is in line with the fact that we only use high quality annotators (§5.3). Further, we manually analyze the 42 annotations of ten randomly sampled annotators: Nine annotators consistently provide high quality annotations, while the other annotator chooses “not a paraphrase” a few times too often (see Appendix C.7 for details). As a result, we assume that most disagreements are due to the inherent plausible label variation of the task (§5.2).
Higher agreement on paraphrase position. Krippendorff’s unitizing $\alpha$ on the highlights is higher than in other areas131313 E.g., 0.41 for hate speech (Carton et al., 2018) or 0.35 for sentiment analysis (Sullivan Jr. et al., 2022). Because of the different tasks these values are not exactly comparable. (see Table 6). We also calculate the “Intersection-over-union” between the highlighted words (i.e., Jaccard Index), a common and interpretable evaluation measure for annotator highlights (Herrewijnen et al., 2024; Mendez Guzman et al., 2022; Mathew et al., 2021; Malik et al., 2021). It seems that while annotations vary on whether there is a paraphrase or not, they agree frequently on the position of the possible paraphrase. On average, at least 50% of the highlighted words are the same between annotations.141414 100% overlap in highlighting is uncommon. DeYoung et al. (2020) consider two highlights a match if Jaccard is greater than 50%. Agreement is higher on the host utterance, because on average the host utterance is shorter than the guest utterance (33 $<$ 85 words).
@@
Preds Shortened Examples
T G D
G: He was the most famous guy in the world of sports… H: The most famous Italian…
G: A lot of them were the Bay Area influx that came up and bought homes to flip. You know what flipping is, right? H: Mm-hmm. Buying a house, improving it, selling it out of profit.
@@ 6 Modeling
Token Classifier. Similar to Wang et al. (2022a), we fine-tune a large DeBERTa model151515microsoft/deberta-v3-large (He et al., 2020) on token classification to highlight the paraphrase positions (for hyperparameters, see App. D.2). We train two models: using all 3,896 training annotations (“ALL” in Table 8) and using the majority aggregated training annotations over the 420 unique (guest, host) training pairs (“AGGREGATED” in Table 8). We consider a model to have predicted a paraphrase for a pair if at least one token is highlighted with softmax probability $\geq 0.5$ in both texts. For each model, we average performances over three seeds.
In-Context Learning. We further prompt the following generative models (see URLs in App. D.1) to both classify and highlight the position of paraphrases: Llama 2 7B and 70B (Touvron et al., 2023), Vicuna 7B (Zheng et al., 2023), Mistral 7B Instruct v0.2 (Jiang et al., 2023), Openchat 3.5 (Wang et al., 2023), Gemma 7B (Team et al., 2024), Mixtral 8x7B Instruct v0.1 (Jiang et al., 2024) and GPT-4161616API calls where performed using the “gpt-4” model id in March 2024. (Achiam et al., 2023). We design the prompt to be as close as possible to the annotator training using a few-shot setup (Brown et al., 2020; Zhao et al., 2021) with all 8 examples shown during annotator training. We also provide explanations in the prompt (Wei et al., 2022b; Ye and Durrett, 2022) and use self-consistency by prompting the models 10 (GPT-4 and Llama 70B: 3) times (Wang et al., 2022b). For the prompt and further hyperparameter settings see App. D.1.
Shortened Example
G: … then he goes on and references and makes mentionof Rudy Giuliani three times in this conversation
H: And Rudy Giuliani was a private lawyer not a government official, so why is he coming up so much in this conversation between two world leaders?
Table 10: Highlighting Differences. We show examples of highlights made by DeBERTa, GPT-4 and human highlights. Lower intensity means less human annotators selected the word. While GPT-4 struggles with providing highlights at all (c.f. extraction error in Table 8), DeBERTa highlights tend to be too sparse (just “Rudy Giuliani”, “coming” and “conversation” in the host utterance). Here, we highlight words, when the softmax probability is >0.44absent0.44>0.44> 0.44181818We tried a few different thresholds $>0.40$ with $0.44$ getting the biggest gain in the Jaccard Index on the test set. instead of $\geq 0.5$. On the complete test set, this also increases the mean Jaccard Index (by $0.06$/$0.01$ for guest/host compared to Table 8).
@@
Overall, GPT-4 and Mixtral 8x7B achieve the best results in paraphrase classification. In highlighting, our DeBERTa token classifiers and GPT-4 achieve the best overlap with human annotations. However, due to problems with extracting highlights from model responses (e.g., hallucinations, see App. D.3), our fine-tuned DeBERTa token classifiers are probably the best choice to extract the position of paraphrases. While the DeBERTa AGGREGATED model achieves higher F1 scores, the DeBERTa ALL model has the highest precision out of all models. We provide our best-performing DeBERTa AGGREGATED model (model with seed $202$ and F1 score of $0.76$) on the Hugging Face Hub [^19] https://huggingface.co/AnnaWegmann/Highlight-Paraphrases-in-Dialog and use it in the following error analysis.
Error Analysis. We consider the best-performing classification and highlighting models for error analysis, i.e., GPT-4 and DeBERTa AGGREGATED. We manually analyze a sample of misclassifications, for examples see Table 9. Overall, the classification quality is better for GPT-4. The DeBERTa classifier finds more paraphrases (note that DeBERTa AGGREGATED for seed $202$ has a recall of $0.86$) but also predicts more false positives than GPT-4. For both models, the items with incorrect predictions also show higher human disagreement. The average entropy for human classifications is lower for the correct ($0.45$ for DeBERTa, $0.45$ for GPT-4) than for the incorrect model predictions ($0.59$ for DeBERTa, $0.67$ for GPT-4). DeBERTa highlights shorter spans of text (on average $6.6$/$6.2$, compared to $16.7$/$10.9$ for GPT-4 for guest/host respectively), while GPT-4 usually highlights complete (sub-)sentences. GPT-4 highlights are largely of good quality, however they often can not be extracted (see App. D.3). The DeBERTa highlights can seem “chopped up” and missing key information (e.g., the original host highlights in Table 10 are just “Rudy Giuliani”, “coming” and “conversation”). We recommend performing a classification of an utterance pairs as a paraphrase when there exist softmax probabilities $\geq 0.5$ for both guest and host utterance, but then selecting the highlights also based on softmax probabilities lower than $0.5$. Alternatively, the best DeBERTa ALL model [^20] https://huggingface.co/AnnaWegmann/Highlight-Paraphrases-in-Dialog-ALL provides fewer but seemingly more consistent highlights (see Appendix D.3). One possible reason for this could be that DeBERTa ALL was trained on individual highlights provided by single annotators, rather than on aggregated highlights.
@@ 7 Conclusion
A majority of work on paraphrases in NLP has looked at the semantic equivalence of sentence pairs in context-independent settings. However, the human dialog setting is highly contextual and typical methods fall short. We provide an operationalization of context-dependent paraphrases and an up-scalable hands-on training for annotators. We demonstrate the annotation approach by providing 5,581 annotations on a set of 600 turn pairs from news interviews. Next to paraphrase classifications, we also provide annotations for paraphrase positions in utterances. In-context learning and token classification both show promising results on our dataset. With this work, we contribute to the automatic detection of paraphrases in dialog. We hope that this will benefit both NLP researchers in the creation of LLMs and social science researchers in analyzing paraphrasing in human-to-human or human-to-computer dialogues on a larger scale.
@@
Even though the number of our unique text pairs is relatively small, we release a high number of high quality annotations per text pair (5,581 annotations on 600 text pairs). Releasing more annotations on fewer “items” (here: text pairs), has increasingly been more common in NLP (Nie et al., 2020; Sap et al., 2022). Further, big datasets become less necessary with better generative models: Using only eight paraphrases pairs in our prompt already led to promising results. We further use the full 3,896 annotations from the training set to train a token classifier showing competitive results with the open generative models. However, the token classifier and other potential fine-tuning approaches would probably profit from a bigger dataset.
@@
Our data creation process was not aimed at scalability. While our developed annotator training procedure can easily be scaled to a larger group of crowd-workers, we manually selected text pairs for annotation. Future work could scale this by skipping manual selection and accepting a more imbalanced dataset or using our trained classifiers as a heuristic to identify likely paraphrases.
@@
In this work we collect a high number of human annotations per item and highlight the plausible label variation in our dataset. However, we use hard instead of soft-evaluation approaches (Uma et al., 2021) for the computational models. We do this because, among others, extracting label distributions for in-context learning is challenging (Hu and Levy, 2023; Lee et al., 2023). We leave the development of a soft evaluation approach to future work but want to highlight the potential of our dataset here: The high number of annotations per item enables the modeling of classifications and text highlights as distributions, similar to Zhang and de Marneffe (2021). Further, our dataset provides anonymized unique ids for all annotators and enables modeling of different perspectives, e.g., with similar methods to Sachdeva et al. (2022) and Deng et al. (2023).
@@
We confirm to have read and that we abide by the ACL Code of Ethics. Beside the mentioned ethical considerations, we do not foresee immediate risks of our work.
Acknowledgements
We thank the anonymous ARR reviewers for their constructive comments. Further, we thank the NLP Group at Utrecht University and, specifically, Elize Herrewijnen, Massimo Poesio, Kees van Deemter, Yupei Du, Qixiang Fang, Melody Sepahpour-Fard, Shane Kaszefski Yaschuk, Pablo Mosteiro, and Albert Gatt, for, among others, feedback on writing and presentation, discussions on annotator disagreement and testing multiple iterations of our annotation scheme. We thank Charlotte Vaaßen, Martin Wegmann and Hella Winkler for feedback on our annotation scheme. We thank Barbara Bziuk for feedback on presentation. This research was supported by the “Digital Society - The Informed Citizen” research programme, which is (partly) financed by the Dutch Research Council (NWO), project 410.19.007.
@@ B.1 Preprocessing
1. Filtering for 2-person interviews. We filter 49,420 NPR and 414,176 CNN interviews from Zhu et al. (2021) for 2-person interviews only. This can be challenging: In the speaker list, authors sometimes have non-unique identifiers (e.g., ‘STEVE PROFFITT’, ‘PROFFITT’ or ‘S. PROFFITT’ refer to the same speaker). If one author identifier string is contained in the other we assume them to be the same speaker.222222 There might be other cases where different string identifiers in the dataset refer to the same speaker although they are not substrings of the other (e.g., ‘S. PROFFITT’ and ‘STEVE PROFFITT’). For a randomly sampled selection of 44 interviews that were identified as more than 2 person interviews, 12 contained errors in the matching. 2/12 were the result of typos and 10/12 were the result of additions to the name like “(voice-over)” or “(on camera)”. We generally assume the first speaker to be the host. We remove 538 NPR and 1,917 CNN interviews because the identifier of the second speaker includes the keywords “host” or “anchor” — thus contradicting our assumption. This leaves 14,000 NPR and 50,301 CNN 2-person interviews.
@@
2. Removing first and last turns of an interview. The first turns in our 2-person interviews are usually (reactions to) welcoming addresses and acknowledgments by host and guest [^23] For example, “I’m Farai Chideya.” “Welcome.” “Thank you.” , while the last often contain goodbyes or acknowledgments242424 For example the last 3 turns in the considered NPR-4 interview: “Well, Dr. Hader. Thanks for the information.”, “Well, thank you for helping share that information […]”, “Well, thanks again. Dr. Shannon Hader […]”. We remove the first two and the last two (guest, host)-pairs. This step removes 2,409 NPR and 26,419 CNN interviews because they are fewer than 5-turns long. For the remaining interviews, this removes 34,773 NPR and 71,646 CNN (guest, host)-pairs.
@@
3. Removing short and long utterances. We further remove short guest utterances of 1–2 words as they leave not much to paraphrase.252525 We manually looked at a random sample of $0.3\%\approx 48$ such pairs. The 1-2 token guest utterances are mostly (40/48) assertions of reception by the guest (e.g., “Yes.”, “Exactly. Exactly.”, “That’s right”). Some are signals of protest (4/48) (e.g., “Hey, man.”, “Yes, but…”, “Hold on.”). None of them were reproduced by the host in the next turn. 3,540 NPR and 12,675 CNN pairs are removed like this. We also remove pairs where the host utterance consists of only 1–2 words.262626 We manually looked at a random sample of $0.3\%\approx 37$ such pairs. The 1–2 tokens host utterances are mostly (28/37) assertions of reception by the host (e.g., “Yeah.”, “Yes.”, “Sure.”, “Right.”, “Right. Right.”, “Ah, okay.”). Some are requests for elaboration (5/37) (e.g., “How so?”, “Like?”, “Four?”) or reactions (3/37) (e.g., “Wow!”, “Oh, interesting.”). Only one example “Four?” was reproducing content in the form of a repetition. . 2,940 NPR and 11,389 CNN pairs are removed like this. We also remove pairs where guest or host utterance consist of more than 200 words.272727 200 is the practical limit for the number of words for the chosen type of question (i.e., ‘Highlight” Question) in the used survey hosting platform (i.e., Qualtrics). It also limits annotation time per question. Overall, this leaves 148,522 (guest, host)-pairs in 34,419 interviews for potential annotation, see Table 11.